### v2

Abstract
A self-consistent model is developed to explain the anomalously large enrichment of nickel isotopes observed in ablation plumes from ultrafast laser irradiation of solid surfaces. The model is based on the spontaneous creation of a magnetic centrifuge in the ablation plume and the associated cyclotron rotation of plasma ions with rotation rates on the order of $10^{9}$ radians per second. Mass separation occurs around the radial coordinate of cylindrical symmetry with longitudinal axis normal to the ablating surface. A Gaussian shaped radial magnetic field distribution is extracted for Ni isotopes with peak field of 53 Megagauss and average field of 20 Megagauss. In addition to cyclotron rotation of ions, a rigid rotor model is also presented that is associated with the hydrodynamic rotation of the entire plasma and is shown to be of little consequence for the isotope enrichment. Cyclotron rotations dominate the process. A profound resonance of enrichment is observed for specific cyclotron charge states and is shown to be associated with Ion Bernstein Wave (IBW) coupling that provides electrostatic acceleration to these resonant orbits.
1 Introduction
In a paper published in 1999 we reported unusually strong isotope enrichment in the normal ejection direction from ultrafast femtosecond laser ablation pulses using laser power densities in the $10^{13}-10^{15}$ W/cm2 Pronko et al. (1999). It was postulated, at that time, based on the angular distribution of those enrichment and concentration distributions of deposited material, that an intense Megagauss magnetic field was responsible for the observed effects. Since then, a number of papers have appeared in the literature reporting on experimental observations of such fields Kahaly et al. (2009); Pisarczyk et al. (2015); Sandhu et al. (2002, 2001) . In view of these developments and the ongoing interest in such magnetic fields, we are revisiting our earlier work and providing important additional data that was obtained during that period, and that has hitherto not been published but which bear directly on the subjectVan Rompay (2003). At the time of that earlier 1999 paper, we were criticized for proposing a magnetic field effect for our observed results, with alternative explanations being proposed using hydrodynamic separation as being responsible for the observed phenomena Gupta and Naik (2001). This criticism was leveled at our work, at that time, out of an incomplete understanding of the characteristics of ultrafast ablation plumes and the various time-dependent processes that occur during its temporal development. In addition, thermodynamic arguments that are relevant to nanosecond laser ablation were being referenced and used in those criticisms. The difference between femtosecond and nanosecond laser ablation pulses is now known to be associated with the fact that femtosecond ablation plumes are characterized by, among other things, a Coulomb explosion emanating out of a solid density plasma VanRompay et al. (1998) and not by the thermodynamic plasma formation associated with nanosecond ablation plasmas.
In the present paper, we will present data showing that the elemental and isotopic spatial separations occurring in ultrafast laser ablation plumes can be consistently described by a magnetic centrifuge process occurring in the self-generated magnetic fields of these ultrafast ablation plasmas. Further we shall show that the magnetic centrifuge can be described by specific Larmor orbits of ions within the plasma with possible assistance from an ExB electric field driven bulk plasma rotation. Details concerning resonant separation enhancement of specific charge states by internal wave excitation will also be presented and discussed.
1.1 Ablation Dynamics
The generation of a plasma from a solid surface by a laser pulse involves the absorption of photon energy by the electrons during the time of the laser pulse and the transfer of that energy to the atoms and ions in the material structure. This transfer of energy to the atoms ultimately results in an ablation plume emanating from the surface. The process of that energy transfer is specific to the laser pulse characteristics and the internal dynamics of the plasma formation. It is important to recognize here that the ablation plume, which is emitted normal to the target surface, will have both spacial and temporal characteristics. We examine the spacial distribution of ions in this report but need to recognize that the plume develops not only with ions, but also with neutrals and nano-particles. This occurs along a specific time line where the pulse absorption is characterized initially by a solid density plasma with fast and slow ions followed by neutrals, and subsequently followed by a burst of nano-particles Albert et al. (2003). These events are observed to have individual time horizons and can, in principle, be separated and studied using mechanical or electrostatic separation mechanisms. Harvesting isotopic ions lends itself to some form of electrostatic separation which will be discussed at a later time.
In regard to the initial plasma state, it is considered axiomatic that a plasma will self-organize after being formed as it relaxes towards a quasi- equilibrium transitional state before eventually dispersing Boyd and Sanderson (2003); Bittencourt (2013). The specific ways this happens will depend on factors such as the electro-magnetic field details that are present and the characteristics of the plasma itself. In the case of high intensity ultrafast laser pulses, as used in this work, the pulse time scales are femtoseconds and intensities on the order of $10^{14}$ to $10^{15}$ watts/cm2. Due to the short time scales and the rapid absorption of energy by the lighter electrons, a coulomb explosion occurs at the surface whereby hot electrons which are ejected outward create a charge imbalance leaving a positive surface layer of ions that are accelerated outward giving them energies in the keV range (see Appendix 8.1) VanRompay et al. (1998). This explosion and transfer of energy happens over a time scale on the order of a few picoseconds or less. During this time interval, electrons also circulate above the target surface and along the low density plasma sheath, acting together with internal thermoelectric currents to form toroidal and longitudinal magnetic fields within the plasma.
The expansion of the initial solid density plasma can be considered to evolve as a one-dimensional (1-D) process, moving as an expanding plug along the normal to the surface of the ablating material Puell et al. (1970). The conversion to a three-dimensional (3-D) expansion occurs after a distance that is on the order of or somewhat greater than the laser beam spot size Stevefelt and Collins (1991). Within the one-dimensional region of the expanding plasma, both toroidal and longitudinal magnetic fields of very high intensity have been observed and measured Kahaly et al. (2009); Pisarczyk et al. (2015); Sandhu et al. (2002, 2001).
The toroidal, or azimuthal, field is considered to form first from rapid outflow of hot electrons under the influence of such mechanisms as the Biermann battery effect. Subsequently a longitudinal field along the one-dimensional z axis forms due to a dynamo effect from a radially expanding, slightly rotating plume that will naturally convert some of a strong toroidal field $B_{\theta}$ into an axial (longitudinal) field $B_{z}$. This occurs primarily through the MHD induction term $\nabla\times(\mathbf{v}\times\mathbf{B})$ together with turbulence such as rotor rotation, tangential surface loops, and non-symmetric three dimensional charge expansion to establish, within picoseconds, a well defined longitudinal field Briand et al. (1985). In addition, the toroidal field will continue to act producing a magnetic pinch on the ionic component of the plasma Yutong et al. (2000); Jackson (2021). This pinching effect will work to hold the plasma together within the one-dimensional zone. The longitudinal magnetic field induces a rotational motion to the expanding plasma and to the energetic particles within, occurring through transverse velocity components and drift rotation, similar to the way it occurs in a conventional plasma centrifuge device Krishnan et al. (1981). The combined effect of these two phenomena, pinching and rotation, can be expected to induce a radial separation of ionic masses within the 1-D cylindrical region of the expansion zone. To the extent that the distribution of ions in the cylindrical zone reaches some quasi equilibrium state, it is expected that these rotational processes will achieve a centrifuge type distribution Krishnan et al. (1981). The plasma behavior in magnetic centrifuge devices has been studied in considerable detail. For the case of an electro-magnetically-driven plasma centrifuge, the rotation rate can be one of two distinct angular frequencies, one at the cyclotron frequency, and one at a specific lower rate, proportional to the $\vec{E}\times\vec{B}$ drift rotation. Derivation and discussion of these two rotational mechanisms are provided in references Kim et al. (1987). In conventional rigid-rotor plasma devices, only the lower of the two characteristic rotation rates is usually observed. This most probably is a consequence of the relatively low magnetic field strengths($\approx 5$kG) used in those devices. In contrast, we show below that our data demonstrates ion rotations at cyclotron frequencies, in these ultrafast ablation plasmas, where much higher field strengths are present (10 -50 MG).
When the ions move into the three-dimensional expansion region, they will transform out of cylindrical symmetry and into a divergent conical symmetry characteristic of pulsed-laser plasma expansion. Figure 1 shows an idealized behavior for how these ions could be considered to undergo this geometric transformation process. The figure will also be used to discuss model calculations presented herein. Those ions on the central axis of the cylinder are considered to remain close to the central axis of the expansion cone, whereas those on the extreme edge of the cylindrical radius are expected to expand out to the largest angles of the cone.
1.2 Laser Parameters and Detector-Target Geometry
[…]
Figure 1: Ablation Model for Enrichment Data as a Function of Observation Angle $\theta$
2 Equations and Data
The basic mass-ratio centrifuge equation in cylindrical coordinates, where $m_{1}$ is less than $m_{2}$ is:
$\frac{n_{1}}{n_{2}}=\exp\left[\frac{(m_{1}-m_{2})\omega^{2}r^{2}}{2k_{B}T}\right]$ (1)
which expresses the mass fraction of a Boltzmann distribution within the cylinder geometry. The terms in this Gaussian function are: $n_{1}(r)/n_{2}(r)$ = concentration ratio of mass $m_{1}$ and $m_{2}$ where $m_{1}<m_{2}$, $k_{B}$ the Boltzmann constant, T the plasma ion temperature, $\omega$ the centrifuge angular rotation rate and r the radial extent of the rotating mass.
Figure 1 shows the essential features of the geometry being used for coordinate transition from an initially expanding cylindrical plug of plasma at the point of laser impact to its subsequent expansion into a conical ablation plume.The laser spot diameter is noted as $d_{1}$ whereas $d_{p}$ is the plasma plume diameter. For clarity, the 1-D zone has been expanded; the length is approximately $d_{p}$/2. The coordinate transformation in Fig.1 shows the relationship between the cylindrical geometry of the plasma centrifuge (with mass separation by radius r) and the angular distribution associated with the ablation plasma (with separation mapped into angle $\theta$).
An elementary conformal mapping function is used such that r/rmax = $\theta/\theta_{max}$. Defining a separation factor S(r) = n1(r)/n2(r) and performing the $\theta$ coordinate transformation, the basic centrifuge equation is transformed to angular coordinates (while also being normalized to unity at the cross over from enrichment to depletion at $\theta_{c}$) as shown in Equation 2.
$S(\theta)=\exp\left[\frac{(m_{1}-m_{2})\omega^{2}r_{\max}^{2}}{2k_{B}T\theta_{\max}^{2}}(\theta^{2}-\theta_{c}^{2})\right]$ (2)
2.1 Data
Two different ablation experiments are presented below. An experiment to test the ablative separation of chemical elements has been performed on a Ni/Cu alloy of 45%Ni and 55% Cu. Figure 2 shows the distribution of the Ni/Cu element separation ratios as a function of charge state and angle for the specific angles of 0, 15. 30, and 60 degrees. Charge states +1 through +10 are shown as gray bars, with the highest enrichment labeled. The total enrichment ratio for all ions is plotted as circles. The enrichment ratio is normalized to the element composition ratio. The gray bars represent the contribution of the various charge states from +1 to +10. These data are presented as clusters at the specific angles of 0, 15 ,30, and 60. It is noted that each group has a dominantly enriched charge state along with sub-dominant peaks. These will be discussed further below. Figure 3 shows a similar plot of the $Ni^{58}/Ni^{60}$ isotope ratios, obtained from a pure nickel target, also showing especially strong dominance of particular charge state enrichments. Figure 4 presents ion density distributions as a function of charge state and energy. These ion population data are used to extract a mean kinetic energy per ion to replace the Maxwellian temperature in the ablative plasma. Since the mean kinetic energy of an ideal gas is $(3/2)k_{B}T$ it is appropriate in the present case to replace $2k_{b}T$ with the mean kinetic energy of $1.05$ kV from figure 4 to be used for further computational purposes. Further discussion of Figure 4 is presented in Appendix section 8.1.
Figure 2: Ni/Cu Enrichment Data as a Function of Angle
Figure 3: Ni Isotope Enrichment Data as a Function of Angle
Figure 4: Ion Energy Distributions of $Ni^{58}$ as Function Charge State
3 Experimental
3.1 Analysis of Data
Figure 5 shows a best fit of Equation 2 (i.e. Gaussian curve) to the 4 major data angles from Figures 2 and 3, with the additional requirement that the separation ratio be unity at the cross over point of 22.5 degrees for the Ni/Cu alloy case. This curve is the result of combining the Ni/Cu data with the Ni isotope data, using average composite masses in Eq. 2, providing an overall averaged angular rotation rate of $3.2\times 10^{9}$ rad/sec. The trend of the data points for both isotopes and alloy mimics the centrifuge curve across all angles, with the best fit dominating the enrichment region at angles below the cross-over point from enrichment to depletion in the alloy case. Analysis of these data points will now be done in terms of the specific centrifuge rotations that are expected to be involved.
Figure 5: Best Fit for Normalized Separation Factor using Hybrid of Ion and Isotope Masses
Rearranging Equation 2, the angular rotation rate, at each observational angle, can be expressed, in terms of separation ratios, as:
$\omega(\theta)=\sqrt{\frac{2k_{B}T\theta_{\max}^{2}\left|\ln S(\theta)\right|}{\left|m_{1}-m_{2}\right|r_{\max}^{2}\theta^{2}}}$ (3)
which yields the specific angular rotation rates for each observed angle from Fig. 5.
Angle $\theta$ $\omega$ (rad/s), Ni/Cu $\omega$ (rad/s), Ni58/Ni60
$0^{\circ}$ $4.50\times 10^{9}$ $12.7\times 10^{9}$
$15^{\circ}$ $2.62\times 10^{9}$ $6.51\times 10^{9}$
$30^{\circ}$ $1.43\times 10^{9}$ $2.46\times 10^{9}$
$60^{\circ}$ $1.13\times 10^{9}$ $0.689\times 10^{9}$
Table 1: Angular frequencies $\omega(\theta)$ for Ni/Cu alloy and Ni58/Ni60 isotope data at selected angles.
Equation 3 has a singularity at zero angle; therefore, the values at theta equal to zero are obtained from a best-fit regression of the other 3 data points in each set that is extrapolated back to zero angle. These final equations and curves are presented below in Equations 4 and 5. Figure 6 shows the final fit of the quadratic and Gaussian equations to the rotation rate results and its extrapolation to the zero (center of the plasma column). Those curves are based on the following fitted equations which are used to extract magnetic fields as a function of observation angle $\theta$ (Section 3.2).
$\omega_{\text{isotope}}(\theta)=8.88\times 10^{9}\cdot\exp\left(-\frac{\theta^{2}}{26.79^{2}}\right)\quad\text{[rad/s]}$ (4)
$\omega_{\text{alloy}}(\theta)=1.63\times 10^{6}\cdot\theta^{2}-1.57\times 10^{8}\cdot\theta+4.77\times 10^{9}\quad\text{[rad/s]}$ (5)
(for further discussion of Eq. 5 see appendix section 8.4).
One of the things to notice about Figure 6 is that the isotope data points are systematically above the Ni/Cu data points as predicted by the cyclotron frequency equation. This is consistent with the cyclotron model as expressed by Eq. 3 in as much as the average alloy mass difference of 4.85 is significantly larger than that of the isotopes, which is 2. It predicts that the general or average rotation rate should scale as the inverse square root of the difference in mass for each case. Using a numerical average for each set of rates, one finds that the ratio of the average frequency for the alloy to that of the isotopes is 0.64 compared to the inverse square root of the mass differences which is 0.52 in general agreement with Equation 3.
3.2 Plasma Rotational Dynamics
In ultrafast laser ablation, the intense laser pulse, which itself has a Gaussian spatial profile Diels and Rudolph (2006), initially creates a highly ionized solid density plasma-plug at the point of impact with the target surface. Modeling this ablation zone as a cylinder with radius on the order-of or somewhat-larger-than the beam spot we can expect to have, within that zone: (a) strong electric fields due to charge separation and coulomb explosion, (b) strong, self-generated magnetic fields due to thermal and charge density gradients and hot electron transport, and (c) rapid plasma expansion transitioning from a dense cylindrical plasma to an expanding conical structure.
Within the initial cylinder there may exist two distinct types of rotational motion. One is the so called ExB drift rotation where the plasma undergoes bulk rotation as a whole. This results from the interaction between radial electric fields and axial magnetic fields. The entire plasma column undergoes rigid-body like rotation with angular rotation rate as given by Chen (2016)
$\omega_{d}=\frac{E_{r}}{RB_{z}}$ (6)
where $E_{r}$ is the radial electric field, $B_{z}$ is the axial magnetic field, and R is the rotating column radius. In this case, which is referred to as the rigid rotor model, the plasma is behaving as a hydrodynamic system rather than a particle plasma.
The other is cyclotron (Larmor) rotation involving individual charged particles moving around magnetic field lines. This results from the Lorentz force that compels electrons and ions into helical orbits. Each species rotates with their own cyclotron frequency
$\omega_{c}=\frac{qB}{m}$ (7)
which depends on mass m, charge q, and relevant magnetic field B leading to species separation, whereas the rigid rotor rotation relies on centrifugal acceleration to accomplish separation. Both modes of separation require a longitudinal magnetic field directed along the z axis of the rotating cylinder.
Equation 7 is found to be critical for the present work. It provides a way of extracting magnetic fields based on cyclotron rotation rates extracted from observed mass separation ratios. Shown below in Figure 6 are the results of using Eq. 7 to calculate the magnetic fields experienced by the alloy elements and the isotopes using their extracted rotation rates as shown in Table 1 above. The numerical average taken from all the data points in Fig.6 provides a field value of 19.3 Megagauss [MG] within the ablation cylinder. These data in Fig. 6 are typical field strengths directly observed experimentally and reported in the literature. Kahaly et al. (2009); Sandhu et al. (2001, 2002, 2004); Forestier-Colleoni et al. (2019); Zhang et al. (2000); Pisarczyk et al. (2015, 2017).
Figure 6: Gaussian and quadratic fit to cyclotron rotation rates and magnetic field as function of plasma radius and observation angle.
The Gaussian equation that fits the local magnetic field for the isotope results as a function of the centrifuge cylinder radius is:
$B(r)=53.38\cdot\exp\left(-\frac{r^{2}}{(55.82)^{2}}\right)\quad\text{[MG]},\quad r\text{ in }\mu\text{m}$ (8)
and the quadratic equation used for the Ni/Cu alloy fit to the magnetic field data points is (see Appendix 8.4):
$B(r)=0.002259\,r^{2}-0.4538\,r+28.64\quad\text{[MG]},\quad r\text{ in }\mu\text{m}$ (9)
Given this information about rotation rates, we can estimate the number of rotations an ion experiences in traveling the 125 $\mu$ length of our idealized centrifuge cylinder. We show, in Table 2 below, the distance traveled per rotation and the number of rotations achieved by an 0.2 keV Ni isotope ion traveling the length of the cylinder. It is seen in Figure 4 that a charge state one ion has this representative energy and represents the bulk of the ionic component in the plasma. The results in Table 2 are computed using separation ratios from Figure 3 and rotation rates from Eq.3. In Fig.3 the total enrichment ratios (for all charge states at a given angle) is shown as circles connected by a solid line. The distance traveled per rotation scales as $E^{1/2}$ for the higher energy ions with a commensurate decrease in number of predicted rotations. At 60 degrees (which has the slowest rotation rate), one full rotation would be accomplished by a 3.2 keV ion traveling the 125 microns. This represents a charge state between 4 and 5 according to Figure 9.
[…]
0 $1.270\times 10^{10}$ $2.031\times 10^{-6}$ 61.54
15 $7.809\times 10^{9}$ $3.303\times 10^{-6}$ 37.84
30 $3.491\times 10^{9}$ $7.389\times 10^{-6}$ 16.92
60 $8.265\times 10^{8}$ $3.121\times 10^{-5}$ 4.01
Table 2: Cyclotron rotation rate, distance per rotation, and number of rotations to travel 125 µm for Ni ions at 0.2 keV.
4 Analytical Model
4.1 Currents and Fields
Let us concentrate on the isotope results in this section, leaving the alloy data for additional discussion at a later time. Given that we now have a mathematical model for the magnetic field within the centrifuge cylinder for the isotopes, we can do the following:
1). Extract azimuthal current densities in the ablation plasma cylinder by using Ampere’s law for the current density equation as a function of the measured magnetic fields. 2). Use this current density to extract the total circulating current that sustains the B(r) field and also to establish the electric field driving the magnetization current. 3). Use this driving electric field to extract a rotation rate for the rigid rotor centrifuge model.
Current Density
Starting with the expression for the magnetic field $B_{z}(r)$ and using Ampère’s Law in cylindrical symmetry, and knowing that an azimuthal current density $J_{\theta}(r)$ will generate an axial magnetic field $B_{z}(r)$:
$\oint\vec{B}\cdot d\vec{\ell}=\mu_{0}I_{\text{enc}}\quad\Rightarrow\quad B_{z}(r)\cdot 2\pi r=\mu_{0}\int_{0}^{r}J_{\theta}(r^{\prime})\cdot 2\pi r^{\prime}\,dr^{\prime}$ (10)
(Note:- Theta in these equations is the azimuthal angle around the cylinder axis, not to be confused with the observation angle theta used to extract our data). Differentiating both sides with respect to $r$, results in the local form of Ampère’s Law:
$\frac{dB_{z}(r)}{dr}=\mu_{0}J_{\theta}(r)$ (11)
Using our experimentally determined form of Gaussian magnetic field profile:
$B_{z}(r)=B_{0}\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)$ (12)
and differentiating:
$\frac{dB_{z}(r)}{dr}=-\frac{2rB_{0}}{\sigma^{2}}\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)$ (13)
which when substituted into Ampère’s law gives the azimuthal current density:
$J_{\theta}(r)=-\frac{2rB_{0}}{\mu_{0}\sigma^{2}}\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)$ (14)
Equation 14 represents a differential current element in a ring segment of $2\pi rdr$ encircling the cylinder. Integrating this around the cylinder provides the total azimuthal current enclosed as a function of radius.
$I_{\theta}(r)=\int_{0}^{r}J_{\theta}(r^{\prime})\cdot 2\pi r^{\prime}\,dr^{\prime}$ (15)
Substituting for $J_{\theta}(r^{\prime})$, gives:
$I_{\theta}(r)=-\frac{4\pi B_{0}}{\mu_{0}\sigma^{2}}\int_{0}^{r}r^{\prime 2}\exp\left(-\frac{r^{\prime 2}}{\sigma^{2}}\right)dr^{\prime}$ (16)
The integral evaluates as:
$\int_{0}^{r}r^{\prime 2}\exp\left(-\frac{r^{\prime 2}}{\sigma^{2}}\right)dr^{\prime}=\frac{\sqrt{\pi}}{4}\sigma^{3}\operatorname{erf}\left(\frac{r}{\sigma}\right)-\frac{r}{2}\sigma^{2}\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)$ (17)
where erf is the error function.
Substituting this into Eq. 16 result in the current as a function of radius:
$I_{\theta}(r)=-\frac{\pi^{3/2}B_{0}\sigma}{\mu_{0}}\operatorname{erf}\left(\frac{r}{\sigma}\right)+\frac{2\pi B_{0}r}{\mu_{0}}\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)$ (18)
In the limit of large r, the enclosed current approaches the total current:
$I_{\theta}(\infty)=-\frac{\pi^{3/2}B_{0}\sigma}{\mu_{0}}$ (19)
Figure 7 shows an overlay of the total current, current density, and cyclotron magnetic field as a function of the cylinder radius. It is seen there that the total local current encompassing the cylinder is on the order of $1.3\times 10^{6}$ amperes.
Figure 7: Total current (blue) surrounding the Gaussian magnetic field of the isotopes along with the extracted current density (red). Gray dashed line is the scaled magnetic field added for visual reference.
5 Rigid Rotor Model
In this section, we derive the rigid rotor rotation rate $\omega_{r}(r)$ for a cylindrically symmetric plasma column in which the azimuthal current density $J_{\theta}(r)$ generates a longitudinal magnetic field $B_{z}(r)$. The analysis assumes the plasma is well-conducting, axisymmetric, and governed locally by Ohm’s law in the azimuthal direction. In addition, by virtue of the current density, it is postulated that the electric field driving the azimuthal current also drives the $E\times B$ drift.
We begin with Ohm’s law for the azimuthal electric field as discussed by Breginskii and Bettencourt for a magnetised plasma Braginskii (1965); Bittencourt (2013):
$E_{\theta}(r)=\frac{J_{\theta}(r)}{\sigma_{e}}$ (20)
where $\sigma_{e}$ is the local plasma conductivity, and $J_{\theta}(r)$ is the azimuthal current density. As done previously, the current density is taken as the Gaussian-weighted linear form which was shown appropriate for the isotope field.
$J_{\theta}(r)=-\frac{2B_{0}}{\mu_{0}\sigma^{2}}r\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)$ (21)
where $B_{0}$ is the peak axial magnetic field at $r=0$, $\mu_{0}$ is the vacuum permeability, and $\sigma$ is a characteristic radial scale length for the current and field profiles.
Substituting into Ohm’s law gives:
$E_{\theta}(r)=-\frac{2B_{0}}{\mu_{0}\sigma_{e}\sigma^{2}}r\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)$ (22)
The rigid rotor rotation rate $\omega_{r}(r)$ is defined as:
$\omega_{r}(r)=\frac{E_{\theta}(r)}{rB_{z}(r)}$ (23)
Knowing that the axial magnetic field is also Gaussian:
$B_{z}(r)=B_{0}\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)$ (24)
and substituting $E_{\theta}(r)$ and $B_{z}(r)$ into the expression for $\omega_{r}(r)$:
$\omega_{r}(r)$ $=\frac{-\frac{2B_{0}}{\mu_{0}\sigma_{e}\sigma^{2}}r\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)}{rB_{0}\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)}$ (25)
$=-\frac{2}{\mu_{0}\sigma_{e}\sigma^{2}}$ (26)
It is seen that $r$, $B_{0}$, and exponential terms cancel, leaving a constant:
$\omega_{r}=-\frac{2}{\mu_{0}\sigma_{e}\sigma^{2}}$ (27)
This result indicates that the plasma, for these field conditions, undergoes uniform rigid-body rotation at a fixed frequency, even though both $E_{\theta}(r)$ and $B_{z}(r)$ vary with radius. The rigidity arises due to the specific structure of the azimuthal current profile, which ensures the ratio $E_{\theta}(r)/(rB_{z}(r))$ remains constant.
The rotation rate depends on the plasma conductivity $\sigma_{e}$, determining the plasma’s response to electric fields, the magnetic permeability $\mu_{0}$, and the radial scale length $\sigma$, controlling the spatial distribution of the relevant fields. This model provides a self-consistent foundation for interpreting uniform rotation in ultrafast laser-produced plasmas with strong self-generated magnetic fields. Equation 27 provides the rotation rate needed to consistently sustain the current and field distributions in our self organized conductive plasma. In that regard, the fields are not independent inputs; they are outcomes of the plasma geometry and the conductivity and do not appear explicitly in the rigid rotor rotation equation. These fields are contained implicitly in the electron cyclotron rotation-rate and the electron ion collision frequency.
Since the electrical conductivity is central to these results we provide, in Appendix 8.3, a brief discussion of the conductivity properties in an asymmetric magnetized plasma such as ours. It is shown there that the component of the conductivity tensor perpendicular to the magnetic field takes the form
$\large\sigma_{\perp}=\frac{\sigma_{0}}{1+\left(\frac{\omega_{ce}}{\nu_{ei}}\right)^{2}}$ (28)
where $\sigma_{0}$ is the conductivity in the absence of (or parallel to) the magnetic field, $\omega_{ce}=\frac{eB}{m_{e}}$ is the electron cyclotron frequency, and $\nu_{ei}$ is the electron-ion collision frequency. For sufficiently large values of $\frac{\omega_{ce}}{\nu_{ei}}$ the perpendicular component of the conductivity is diminished due to the action of the magnetic field continuously interrupting the forward velocity of the electrons between collisions with ions. It is shown there that, for our electron temperature and density, the electron ion collision frequency is very much larger than the the electron cyclotron frequency ($\nu_{ei}\sim 10^{18}\,\text{Hz}$, $\omega_{ce}\sim 10^{14}\,\text{rad/s}$). This means the electrons collide with ions well before they complete one full rotation resulting in $\frac{\omega_{ce}}{\nu_{ei}}$ essentially approaching zero. Thus, for our case, the perpendicular conductivity reduces to the field free Spritzer conductivity for a fully ionized plasma which is given by:
$\sigma_{e}\approx 1.96\times 10^{7}\,\frac{T_{e}^{3/2}}{Z\ln\Lambda}\quad[\text{S/m}]$ (29)
where $T_{e}$ is in eV, $Z$ is the ion charge state, and $\ln\Lambda$ is the Coulomb logarithm. The computed value of this conductivity is found to be
$\sigma_{\text{Spritzer}}=1.96\times 10^{9}\,\text{S/m}$
and when used with the gaussian scale length of 56 microns results in a rigid rotor rotation rate of
$\omega_{r}=-2.59\times 10^{5}\,\text{rad/s}$
The negative sign is because the rigid rotation is opposite to the sense of rotation for the cyclotron orbits. Parameters used for these calculations can be found in Appendix 8.3.
5.1 Dominance of Cyclotron Rotation in Isotope Separation
Our analysis of rotation rates of and within the plasma shows that ion cyclotron (Larmor) motion significantly dominates over collective $\vec{E}\times\vec{B}$ rigid rotor drift. Specifically, the measured rigid rotor frequency is four orders of magnitude lower than the corresponding ion cyclotron frequencies for the charge states of interest (see Table 1 above). As a result, while the $\vec{E}\times\vec{B}$ drift contributes to bulk plasma rotation at a global scale, it apparently plays a minimal role in isotope separation dynamics. Instead, separation is primarily governed by individual cyclotron-driven motion, which provides much stronger mass-dependent transport. This observation means that the key mechanism for isotope enrichment in our system arises from differences in Larmor rotation characteristics such as gyro radius and frequency rather than from hydrodynamic bulk processes. The implications are significant: mass-selective separation can occur efficiently provided that localized magnetic field forces are present to couple into the cyclotron trajectories of specific ion species both generally and resonantly.
6 Resonant Isotope Enrichment- Ion Bernstein Waves
The anomalously high isotope entrenchments for specific charge states, as seen in Figure 3 (and to a lesser extent in Figure 2), are suggestive of some kind of resonance enhancement of the enrichment process. This effect can be observed for example in Fig.3 for the specific charge states of Ni-58 at charge states $Z=9,8,6$ , and in particular being more dominant at smaller expansion angles. This behavior is consistent with and is, we propose, caused by resonant coupling between the cyclotron rotating ions and associated ion Bernstein waves (IBWs) that exist within a magnetized plasma Boyd and Sanderson (2003); Bittencourt (2013). IBWs are longitudinal electrostatic waves that propagate perpendicular to the magnetic field and exist in harmonics of the ion cyclotron frequency. These waves can interact resonantly with ions when their wave frequency matches the ion’s cyclotron frequency or one of its harmonics. In our ultrafast laser ablation plasma, the strong, self-generated magnetic field produced in the early expansion phase enables a localized IBW spectrum that overlaps with the higher cyclotron frequencies of high charge state ions. Because these ions are also among the fastest particles in the plasma, they traverse the high-field region rapidly and are preferentially positioned to undergo coherent wave-particle coupling. This coupling leads to selective energy transfer and preferential transport of certain isotopes—such as Ni-58—whose mass-to-charge ratio allows for near-exact resonance with the available IBW modes. The sharpness of these enhancements, and their confinement to specific charge states and angular observation regions, cannot be easily explained by other mechanisms such as thermal diffusion, bulk drift rotation, or other hydrodynamic processes . Instead, they point to a nonlinear, field-mediated resonance mechanism such as IBW coupling, which can selectively amplify isotope populations far beyond what would be expected from normal equilibrium behavior. The presence of IBW-driven enrichment signatures, especially in low-density, high-$Z$ ion populations, thus serves as a compelling indicator of localized, resonance-based separation dynamics acting within the early-time magnetic structure of the plasma.
6.1 Charge-State–Localized Resonance Condition:
In this section we adopt a physical model in which each charge state excites its own local IBW as it traverses the plasma. Each charge state $Z$ couples to an Ion Bernstein Wave (IBW) with a frequency equal to its local cyclotron frequency at its specific resonance radius $r_{Z}$. These local IBW frequencies are not all equal but are proportional to the charge state and the local magnetic field. With the cyclotron frequency being given as:
$\omega_{c,Z}(r_{z})=\frac{Z\cdot e\cdot B(r_{z})}{m_{i}}$ (30)
the expression for IBW matching will be: $\omega_{\text{IBW}}(Z)=\omega_{c,Z}(r_{Z})$.
The dominant resonance in Figure 3 is the +9 charge state at the observation angle of zero degrees where the isotope enrichment reaches a factor of 20 above natural abundance. We take this as the fundamental IBW frequency for the system. Other charge states with integer values of Z will have resonances at sub-integer multiples of this fundamental frequency, and will occur at various radii, based on their location in the gaussian magnetic field. By calibrating to charge state $Z=9$ at $r=0$, a reference point is set up, whereby the fundamental IBW frequency is established for the system along with a set of sub-harmonics. Using Eq.30 and matching the general cyclotron frequency to the fundamental IBW
$\frac{Z\cdot B(r_{z})}{m_{i}}=\frac{9\cdot B(0)}{m_{i}}\quad\Rightarrow\quad B(r_{z})=\frac{9}{Z}B(0)$ (31)
where $B(r_{z})$ is given by our Gaussian field profile. Thus
$\frac{9}{Z}B(0)=B(0)\cdot\exp\left(-\left(\frac{r_{z}}{\sigma_{r}}\right)^{2}\right).$ (32)
Equation 31 defines the magnetic field needed to match the resonance conditions for each charge state and, by combining that with our field expression (Eq. 32), a value of the radius is determined at which a particular Z charge state resonance should appear.
Canceling $B(0)$ on both sides of Eq. 32 and solving for $r$ gives the final expression for the resonance radius at which the charge states will appear:
$r_{z}=\sigma_{r}\cdot\sqrt{\,\left|\ln\left(\frac{9}{Z}\right)\right|},$ (33)
ensuring that lower charge states find resonance farther from the center, where the magnetic field is weaker. $\sigma_{r}$ defines the width of the Gaussian field profile and is discussed next.
6.2 Increasing the Gaussian Width to Match Observed Resonance Positions
In our model, as presented above, the magnetic field profile is described by a Gaussian where $\sigma_{r}$ characterizes the spatial extent of the magnetic field within the plasma. The resonance condition for each charge state is set by matching the ion’s cyclotron frequency to a localized IBW, yielding the resonance radius of Eq. 32. This relation states that the resonance radius scales directly with $\sigma_{r}$. However, when applying this model to experimental data using the initially determined field width (from Eq. 8, $\sigma_{r}=56\,\mu\text{m}$), the predicted resonance positions for lower charge states (e.g., $Z=8,7,6$) fall systematically short of the observed angular locations by a factor of about 3. To bring the calculated resonance positions into agreement with experimental enrichment angles, the Gaussian width must be increased from $\sigma_{r}=56\,\mu\text{m}$ to $\sigma_{r}=182\,\mu\text{m}$ . Physically, this corresponds to a broader magnetic field distribution that drops off more slowly with radius, allowing lower charge states to resonate farther from the center which is where they are actually observed. Increasing $\sigma_{r}$ in this way does not require a higher peak field $B_{0}$, but instead simply reflects a magnetic field that more uniformly fills the centrifuge cylinder providing a broader spatial variation of the magnetic field with radius. This field density relaxes, through dissipative effects in the plasma, to evolve into the more narrow and concentrated quasi-equilibrium field that effects the observed bulk isotope enrichment. See Fig.8 for a comparison of the shapes of these two fields.
Table 3 below shows the results of using Eq. 33 to compute the radius at which the IBW resonances should occur for the various isotope charge states as observed in Figure 3. It is seen there that Equation 33 predicts the location of the resonances for charge states 8 and 6 with very good accuracy. As seen in the data of Figure 3, charge state +9 was dominant in both the zero and 15 degree angles. It has about a factor of 2 less enrichment at 15 compared to the on axis zero degrees. Also, it contains a contribution from charge state +10. These results suggest some complex sub-harmonic mixing is going on at this angle. Calculating the value of the charge state that should resonate at 15 degrees (by inverting Eq.33) results in a theoretical value of Z=8.8 which is reasonably close to the observed charge state +9. It can be concluded that at our 15 degree observation angle the +9 ion cyclotron rate is slightly off the strong resonance as seen to occur at zero degrees.
Z $r_{Z}$ $\theta$ Obs. $\theta$
9 0 0 0 and 15
8 62.50 30.00 30
7 91.29 43.82
6 116.0 55.66 60
[…]
The introduction of these two different, but self similar, magnetic fields implies a certain time evolution of the relaxing magnetic field from its early initial state to the final quasi-equilibrium state. It is understood that the high charge state ions (e.g., $Z=9,8,6$) produced in ultrafast laser ablation are typically those with the highest kinetic energies, originating from the peak of the Coulomb explosion (see Appendix 8.1). These ions travel at significantly higher velocities than the lower charge states and are therefore the first to traverse the magnetic field region immediately following plasma formation. This early-time field, generated through various initiating mechanisms and dynamo processes, is characterized by a broader and more relatively uniform field distribution. It is during this initial phase that coherent resonant coupling with ion Bernstein waves (IBWs) is dominating with the high-$Z$ ions. These interactions leave distinct enrichment signatures that appear anomalous when compared to the behavior of the bulk plasma. As the plasma evolves, the initially broad field relaxes into our Gaussian distribution with $\sigma_{r}=56\,\mu\mathrm{m}$ but retaining its peak value of $B_{0}=53.38\,\mathrm{MG}$ at the core of the cylinder. This temporal sequence supports a model in which early high-field interactions influence the motion of fast ions, while the later-stage, lower-field configuration governs the mass-selective separation dynamics of the slower, more abundant, low-$Z$ species (See Figs. 4 and 9). In both cases the peak field at the core of the cylinder remains the same, with only the Gaussian width $\sigma_{r}$ changing.
The persistence and relaxation time of the two field types can be estimated by examining the ion transit times of the various fast and slow ions through the magnetic centrifuge cylinder. The fastest ions (e.g., charge state $Z=9$ to $6$) traverse the early, wide, high density Gaussian field. These high-$Z$ ions are created at the peak of the Coulomb explosion and have the highest energy. They move rapidly (tens of keV), reaching the end of the 125 micron plasma column in the $500$ to $800$ picoseconds range. These higher velocity ions are the ones being supported by ion Bernstein wave (IBW) resonances. The slowest ions (e.g., $Z=1$, representing the bulk plasma) traverse the narrower, equilibrium Gaussian magnetic field.
Figure 8: Required magnetic field (dashed line) needed to satisfy the IBW enrichment resonances of high charge state (i.e. high velocity) ions compared to the field profile (solid line) used for the low velocity enrichment of the isotopes. Both curves are Gaussian but with differing width. (High velocity-182 micron, low velocity-56 micron).
These ions are more abundant and have lower energies ($\approx 0.2$ keV), resulting in slower transit times of a few nanoseconds. They experience the relaxed magnetic field structure, which governs the bulk isotope separation. This sets the upper and lower bounds for the magnetic field creation and collapse time scales. Using the energy vs charge state data of Fig.10 (in Appendix section 8.1) for velocity information, the early magnetic field with wide gaussian must form on a very short time scale, significantly shorter than 537 picoseconds, which is the 125 micron transit time of the fastest +9 ion at 16.3 keV. This early field needs to last for a time of 835 picoseconds which is the transit time of the charge state +6 ion at 6 keV which is observed resonating at 60 degrees and is the slowest of the resonating ions. The final quasi-stationary field of the 56 micron Gaussian width needs to last for a time that is at least 4.85 nanoseconds to clear the transit time of the slowest low energy +1 ions at approximately 0.2 keV..
While high charge state ions (e.g., $Z=9,8,6$) exhibit strong and selective resonant enhancements in our experimental data, it is important to recognize that these ions represent a very small fraction of the total plasma population being three to four orders of magnitude lower in density than the dominant low charge states ($Z=1$ and $Z=2$). Consequently, although the high-$Z$ species provide clear evidence of coherent wave-particle interactions, such as coupling to ion Bernstein waves (IBWs), they are unlikely to drive the bulk of the observed isotope separation. Instead, the primary separation process is governed by the dynamics of the much more abundant low charge state ions, which respond to the overall plasma structure. In this context, the high-$Z$ resonant enhancements serve primarily as diagnostic tracers of underlying plasma wave activity and magnetic field structure during the early stages of plasma evolution. They provide key information about the plasma’s resonant behavior without being directly responsible for the majority of mass separation observed at later quasi-equilibrium stages. It is instructive to think about how these resonances could be brought down to the low Z ions in order to enhance energy coupling to them and the bulk of the mass enrichment process. It would require much higher magnetic fields to accomplish this along with possibly using very low mass elements such as lithium, beryllium, or boron since cyclotron frequencies scale linearly with B and with mass as 1/m.
7 Conclusion
An in-depth examination of ion dynamics associated with ultrafast laser ablation plumes, from $5.2\times 10^{15}$ $W/cm^{2}$ pulses, emanating from nickel and a nickel/copper alloy surface provide detailed information about self-generated magnetic centrifuge fields in these ultrafast ablation plasmas. We examined, in particular, the effects of those fields on the radial mass enrichments around an axis normal to the ablating surface. It is shown that these energetic ablative ions, being subjected to a cylindrically symmetric longitudinal magnetic field which is oriented normal to the ablating surface, respond to a centrifuge mechanism that drives the radial distributions. Magnetic fields with peak values of 53 MG and average fields of 20 MG are implied from isotope enrichment data with local azimuthal electric currents on the order of $1.3\times 10^{6}$ amperes driving these fields.
The analysis of isotope enrichment and our subsequent findings proceeded as follows: (1). Obtain mass separation ratios by casting the standard centrifuge equation into the initial geometry of an ablating cylindrical model that, after a short distance, transitions to an expanding cone. Apply this model of an expanding cone to find angular distributions of observed mass enrichments in terms of separation ratios. (2). These separation ratios are fitted to the centrifuge equation providing an overall averaged angular rotation rate for ions of $3.2\times 10^{9}$ radians/second. (3). The angular data scans are mapped back to the initial cylinder radial coordinate that allows detailed examination of the cyclotron orbits of ions at specific cylinder radii resulting in specific ion rotation rates at those radii. These are then converted to local magnetic field values, through the cyclotron equation, that provides a radial magnetic field distribution within the cylindrical centrifuge. This field distribution then provides a mathematical model that is used to extract current densities and enveloping currents responsible for the magnetic centrifuge mechanism. (4). The magnetic fields and currents are then used to construct the rigid rotor model of plasma rotation resulting in a value of $2.59\times 10^{5}$ radians per second, significantly lower than the spatially averaged cyclotron rates. (5). Anomalous localized charge state enrichments are examined in terms of resonant energy absorption from plasma waves that are identified as Ion Bernstein Waves (IBW). These are found to provide an excellent fit to the observed resonance data. A time dependent radial magnetic field profile is found necessary to properly fit these resonant data points. It is determined that this initial magnetic field lasts for at least 885 picoseconds and thereafter evolves to a stable quasi equilibrium configuration that lasts for as long as 4.85 nanoseconds before dissipating.
8 Appendix
8.1 Coulomb Explosion
As explained earlier, the ablation process under femtosecond laser ablation occurs by a coulomb explosion due to the lighter electrons being stripped from the surface leaving a high concentration of positive ions in a solid density plasma condition. We described this explosion, in a previous report in terms of a parallel capacitor electric field model VanRompay et al. (1998). That model predicts a linear relationship between the charge state and its energy. The data below show that a quadratic behavior is more typical and needs to be analyzed in those terms. See Figure 4 for details where the data points in Figure 9 are taken as the most probable energy at each charge state.
Figure 9: Mean ion energy as function of charge state shows super liner (quadratic) coulomb explosion. See Fig.4 for details.
These points can be fit nicely to a quadratic equation passing near the origin.
$E(Z)=0.244\,Z^{2}-0.399\,Z+0.094\quad\text{[keV]}$ (34)
This result can be interpreted as higher charge states experiencing stronger net forces during the coulomb explosion suggesting that higher $Z$ ions originate from deeper within the initially formed solid density plasma, where the local positive potential is strongest, thus experiencing the largest forces and gain the most energy. The non-linearity reflects a collective effect not just from its own charge, but through interaction with the entire charge cloud.
In a Coulomb explosion, each ion is repelled by the net positive charge of its neighboring ions. The resulting force on an individual ion scales approximately as:
$F\propto Z\cdot E_{\text{local}}$
where $Z$ is the ion charge state and $E_{\text{local}}$ is the local electric field generated by the surrounding ion cloud. Importantly, this field is not constant—it reflects the cumulative contribution of neighboring ions, which also depend on their respective charge states $Z_{j}$.
The energy gain of an ion depends on both its own charge and the electric potential it experiences. This can be approximated as:
$E_{i}\propto Z\cdot\Phi_{\text{Coulomb}}\propto Z\cdot\left(\sum_{j}\frac{Z_{j}}{r_{ij}}\right)$
where $r_{ij}$ is the distance between ion $i$ and neighboring ion $j$.
In a dense central region of the exploding plasma, the potential $\Phi_{\text{Coulomb}}$ scales linearly with the total enclosed charge, leading to a quadratic type dependence of energy on charge state:
$E_{i}\propto Z^{2}$
This relationship aligns with experimental observations showing that higher-$Z$ ions emerge with disproportionately larger kinetic energies. It reflects the collective nature of the explosion, where the energy is not solely a function of individual charge, but of coherent interaction with the entire charge distribution.
8.2 Collimation Restrictions
The sector analyzer has a 2 mm acceptance aperture at its input Van Rompay (2003); VanRompay et al. (2000). The angular resolution $\Delta\theta$ of this ion detector is the half-angle of the acceptance cone subtended by the detector aperture as seen from the target:
$\Delta\theta=\tan^{-1}\left(\frac{r_{\text{aperture}}}{d}\right)$
Where: $r_{\text{aperture}}$= radius of the detector aperture and $d$ is the distance from the target to the detector
$r_{\text{aperture}}=$1.0\text{\times}{10}^{-3}\text{\,}\mathrm{m}$\quad\text{and}\quad d=$1.1\text{\,}\mathrm{m}$\$
[…]
Using the angle to radius conversion from Fig.1
$\frac{r}{r_{\text{max}}}=\frac{\theta}{\theta_{\text{max}}}\quad\Rightarrow\quad r=r_{\text{max}}\cdot\frac{\theta}{\theta_{\text{max}}}$
$r=1.25\times 10^{-4}\cdot\frac{0.052}{60}=1.083\times 10^{-7}m\$
[…]
8.3 Conductivity Perpendicular to B
Appendix: Perpendicular Conductivity in a Magnetized Plasma
Begin with the steady-state force balance equation for electrons in a magnetized, collisional plasma Boyd and Sanderson (2003); Bittencourt (2013):
$e\vec{E}=-e\vec{v}_{e}\times\vec{B}-m_{e}\nu_{ei}\vec{v}_{e}$
This equation reflects the balance between the electric force, the Lorentz force from the magnetic field, and the collisional drag acting on electrons due to scattering from ions. Assuming linear response, this equation determines the steady-state electron drift velocity $\vec{v}_{e}$ in terms of the applied electric field $\vec{E}$.
To solve this equation, we choose a coordinate system where the magnetic field is aligned along the $\hat{z}-$axis, i.e., $\vec{B}=B\hat{z}$, and write the drift velocity as $\vec{v}_{e}=(v_{x},v_{y},v_{z})$. The cross product $\vec{v}_{e}\times\vec{B}$ then becomes $(v_{y}B,-v_{x}B,0)$, and the collisional drag term is simply $\nu_{ei}\vec{v}_{e}$, where $\nu_{ei}$ is the electron-ion collision frequency.
Substituting into the force balance and equating components yields:
$-E_{x}$ $=\nu_{ei}v_{x}+\omega_{ce}v_{y}$
$-E_{y}$ $=\nu_{ei}v_{y}-\omega_{ce}v_{x}$
$-E_{z}$ $=\nu_{ei}v_{z}$
where $\omega_{ce}=eB/m_{e}$ is the electron cyclotron frequency. Solving this system for $\vec{v}_{e}$ in terms of $\vec{E}$ leads to the expressions:
$v_{x}$ $=\frac{-\nu_{ei}E_{x}-\omega_{ce}E_{y}}{\nu_{ei}^{2}+\omega_{ce}^{2}}$
$v_{y}$ $=\frac{-\nu_{ei}E_{y}+\omega_{ce}E_{x}}{\nu_{ei}^{2}+\omega_{ce}^{2}}$
$v_{z}$ $=-\frac{E_{z}}{\nu_{ei}}$
From these, the current density $\vec{J}=-en_{e}\vec{v}_{e}$ is computed, yielding a linear relation between $\vec{J}$ and $\vec{E}$ through a conductivity tensor:
$\vec{J}=\boldsymbol{\sigma}\cdot\vec{E}$
The resulting conductivity tensor takes the form:
$\boldsymbol{\sigma}=\sigma_{0}\begin{bmatrix}\frac{1}{1+b^{2}}&-\frac{b}{1+b^{2}}&0\\ \frac{b}{1+b^{2}}&\frac{1}{1+b^{2}}&0\\ 0&0&1\end{bmatrix}\quad\text{with}\quad b=\frac{\omega_{ce}}{\nu_{ei}}$
Here, $\sigma_{0}$ is the classical Spitzer conductivity (for $\vec{B}=0$), and $b$ is the dimensionless magnetization parameter.
The diagonal entries of the tensor describe the conductivity parallel and perpendicular to the magnetic field. The off-diagonal elements represent the Hall conductivity, which accounts for the rotation of the current vector due to the Lorentz force. Specifically, the conductivity perpendicular to the magnetic field is given by:
$\sigma_{\perp}=\frac{\sigma_{0}}{1+\left(\frac{\omega_{ce}}{\nu_{ei}}\right)^{2}}$
This result shows that the presence of a magnetic field suppresses the ability of electrons to carry current across field lines. In the limit of strong magnetization ($\omega_{ce}\gg\nu_{ei}$), the perpendicular conductivity becomes significantly smaller than the un-magnetized case. In contrast, the conductivity parallel to the field remains unaffected and equals $\sigma_{0}$. It is seen that the components of the tensor are controlled by the electron cyclotron frequency and the electron-ion collision rate. These are primarily controlled by the electron density, electron plasma temperature, and magnetic field strength.
Table 4: Key Assumptions for Conductivity Calculation in 125 µm Plasma Cylinder
Parameter Assumed Value / Condition
[…]
Plasma species Fully ionized Ni or Ni/Cu alloy
Charge state $Z=1$
Electron temperature $T_{e}=100\,\text{eV}$
Electron density $n_{e}\sim 10^{24}-10^{26}\,\text{m}^{-3}$
Coulomb logarithm $\ln\Lambda=10$
Magnetic field strength $B(\theta)\leq 5338\,\text{T}$ (Gaussian profile)
Plasma regime Fully ionized, weakly coupled, collisional
Conductivity model Spitzer conductivity with perpendicular suppression
Conductivity formula $\sigma_{\perp}=\frac{1.96\times 10^{7}\,T_{e}^{3/2}}{Z\ln\Lambda\left[1+\left(\frac{\omega_{ce}}{\nu_{ei}}\right)^{2}\right]}$
Using the empirical formula for the electron-ion collision frequency Boyd and Sanderson (2003) (with $T_{e}$ in eV, $n_{e}$ in m-3):
$\nu_{ei}\approx 2.91\times 10^{-6}\cdot\frac{n_{e}Z\ln\Lambda}{T_{e}^{3/2}}\quad\text{[Hz]}$
it is found that
$\nu_{ei}\sim 2.8\times 10^{18}\,\text{Hz}$
and more specifically for our isotope gaussian field distribution:
$1.10\times 10^{14}\,\text{rad/s}\leq\omega_{ce}\leq 9.39\times 10^{14}\,\text{rad/s}$}
Based on these electron-ion collision frequencies it is seen that the perpendicular conductivity reduces to the field free Spitzer value.
$\sigma_{0}\approx\frac{1.96\times 10^{7}\,T_{e}^{3/2}}{Z\ln\Lambda}\quad\text{[S/m]}$
8.4 Alloy Gaussian
In both the alloy and the isotopes case it was necessary to get an analytical fit to the first three data points in order to extrapolate to zero frequency to get the magnetic field strength on the central axis (since it is a singularity in equation 3). In the isotopes case the gaussian fit came naturally to the cyclotron frequencies and associated magnetic fields (Figure 6). This is expected since the laser beam profile has a gaussian shape and projects that onto the plasma. Also, the centrifuge equation partitions out mass in a gaussian fashion. This simple Gaussian fitting however was not the case when dealing with the alloy data. It was necessary to first fit a quadratic expression to extrapolate to zero and then subsequently to extract a gaussian. That is what we are doing here. It is found that a gaussian can be arrived at by first subtracting a constant background field from the quadratic expression. This is sown in Figure 10. The constant background is found to be 7.67 MG. The enrichments for the Ni/Cu alloy composition case has a cross over from enrichment to depletion of the light ion at 22.5 degrees whereas the isotope case did not achieve that expected behavior. Why this difference occurred is only partially understood at this point. In Figure 10 a vertical line is shown at that cross over point and it is seen that this is where the extracted gaussian rapidly approaches zero with only the constant background field remaining. This difference between the isotopes and alloy cases is being pointed out as an interesting feature without further elaboration here. Equation 34 is the extracted gaussian expression. The cross over point in the alloy data where it goes from enrichment to depletion of the light ion is at a radius of 48.9 microns (22.5 degrees). This suggests that the enrichment of the heavy element is occurring primarily in the constant background magnetic field. In the decomposition: The extracted Gaussian in Megagauss is:
$B_{\text{Gaussian}}(r)=20.90\cdot\exp\left(-\frac{r^{2}}{(34.52)^{2}}\right)$ (35)
The constant background subtracted from the quadratic is $7.67MG$. The sum of these two yields the green curve in Figure 10.
Figure 10: Alloy quadratc/gaussian decomposition
Acknowledgments
The authors would like to thank the University of Michigan and Gerard Mourou, while director at the Center for Ultrafast Optical Science, for providing support and resources during the experimental phase of this work. The equipment used to collect these data was assembled through a grant from the Air Force Office of Scientific Research.